Magnesium reacts with HCl to produce MgCl2 and hydrogen gas.

How many liters of a 1.85 M HCl solution completely react with 4.00 grams of magnesium?

Mg(s) + 2 HCl(aq) → MgCl2(aq) + H2(g)
1
Expert's answer
2020-11-23T06:55:08-0500

"Mg + 2HCl \to MgCl_2 + H_2"

1 mole of Mg reacts with 2 moles of HCl

24g of Mg = 2(36.5)g of HCl

24g of Mg = 73g of HCl

4g of Mg = x of HCl

x =4×73/24

x = 12.17g

Therefore, 12.17g of Mg is produced =

12.17g/24g/mol = 0.507 mol


To get the volume needed,

Volume = no. of moles/concentration

Volume = 0.507/1.85 = 0.274dm³ = 0.274L.